OpenStudy (anonymous):

Find the volume using cylindrical shells y = 2ex, y = 2e−x, x = 1; about the y-axis

myininaya (myininaya):

I like to start with a rough graph first. Is your equations really y=2ex, y=2e-x,x=1?

myininaya (myininaya):

or did you mean y=2e^x,y=2e^(-x),x=1

OpenStudy (anonymous):

I meant the second one sorry

myininaya (myininaya):